### v2

@@ 2.1 Reconstruction
The first scheme that we identified revolves around the idea of masking some piece of information at any stage during the forward pass of the model, and then employing a decoder to reconstruct the missing data. This pre-training technique was introduced in concurrent works by [9] and [2]. Both works consist of an encoder based on a ViT architecture [22], in which a large portion of the input tokens (resulting from linearly projecting the non-overlapping patches of the image) are masked. After the encoding stage, where masked tokens are substituted with a special token, a decoder is used to reconstruct them. In the Masked Autoencoder (MAE) proposed by [9], the decoder is lighter than the encoder, which results in an asymmetric architecture. The loss function is applied only to the output corresponding to the masked tokens. When applied to a downstream task, the decoder is dropped, only the encoder being used as the backbone due to its strong feature representation capability. We illustrate the reconstruction framework in Figure 2.
@@
Models: $V\!P_{\varphi}$ – the visual projection layer; $E_{\theta}$ – the encoder; $D_{\phi}$ – the decoder.
@@
2:$I_{v},I_{m}\leftarrow\mbox{mask}(P,\alpha,n)$
3:for $i\in\{1,\dots,n\}$ do
4:     if $i\in I_{v}$ then
5:         $H[i]\leftarrow E_{\theta}(V\!P_{\varphi}(p_{i}))$
6:     else if $i\in I_{m}$ then
7:         $H[i]\leftarrow M$
8:$\hat{P}\leftarrow D_{\phi}(H)$
9:$\mathcal{L}\left(\phi,\varphi,\theta,M\right)\leftarrow d\left(\hat{P},P\right)$
10:$\theta\leftarrow\theta-\eta\cdot\frac{\partial\mathcal{L}}{\partial\theta}$
11:$\phi\leftarrow\phi-\eta\cdot\frac{\partial\mathcal{L}}{\partial\phi}$
12:$\varphi\leftarrow\varphi-\eta\cdot\frac{\partial\mathcal{L}}{\partial\varphi}$
@@
We formally present the reconstruction-based training strategy in Algorithm 1. The first step of the algorithm splits the input images into non-overlapping patches, resulting in a set $P$ that contains patches of the same size, namely of $h\times w$ pixels. The second step applies the masking operation, which can differ from one method to another. The masking operation indicates which patches should be kept and which should be masked. The indexes of the visible and masked patches are stored in $I_{v}$ and $I_{m}$, respectively. In the next steps, the masked patches are usually dropped, and only the visible patches are processed by a visual projection layer ($V\!P_{\varphi})$, followed by an encoder ($E_{\theta}$) that extracts a latent representation. Before the decoding step, the previously masked patches are replaced by a learnable representation ($M$), which resides in the latent space of the encoder. These transformations correspond to steps 3-7 in Algorithm 1.
Using the sequence formed by concatenating the masked representations with those from the encoder, the decoder ($D_{\phi}$) reconstructs the patches as depicted in step 8 of Algorithm 1. Finally, in steps 9-13 of Algorithm 1, a distance function is optimized by updating the encoder, the decoder, the projection layer, and the learnable masked representation.
@@ 2.2 Contrastive
Figure 4: A multi-level classification of Masked Image Modeling papers into various categories, based on the research directions studied in the respective papers. References are clickable links to papers.
@@ 3.1 Masking Strategy
A number of studies used low-level information to improve the masking strategy. [52] formulated the masking strategy as as curriculum learning problem by gradually increasing the masking ratio during training. Additionally, they introduced a patch selection strategy that emphasizes masking patches with higher gradient magnitudes. The masking strategy proposed by [54] is inspired by the color noise used in image processing. Specifically, their approach introduces four filters that operate on the uniform noise typically generated for random masking. This strategy avoids the extra computational overhead often required by other informed masking methods while still outperforming standard random masking. Rather than imposing a masking policy, several studies tried to learn the optimal masking strategy. SemMAE [42] deployed an additional stage before masking. This stage, called semantic part learning, is responsible for learning attention maps that correspond to meaningful semantic components in the image. The training of this stage is performed by embedding the class token provided by a ViT encoder into part embeddings. The resulting embeddings together with the patch embeddings provided by the same encoder are used in an attention layer. Finally, the obtained attention maps are processed by a decoder to reconstruct the original image. In the second stage, the attention maps are used for part segmentation. The masking varies from masking patches in each part to masking entire parts randomly. In a similar fashion, [41] proposed a framework in which the masking strategy is learned. A teacher network, seeing the intact image, generates a mask that is used to pre-train the student. The objective is to reconstruct the feature representation of the teacher for the masked tokens, as well as the class token [CLS] that is used in generating the mask. The teacher’s parameters are updated using an exponential moving average of the student’s weights. The experiments demonstrated the superiority of the method over other masking strategies. Instead of relying on an additional teacher network, [47] developed a masking strategy based on the direct feedback provided by the model being pre-trained. The strategy computes the reconstruction loss for each patch and then selectively masks the patches that are more challenging to reconstruct, as indicated by their higher loss values. This approach ensures that the model focuses on the most difficult aspects of the data during training. Similarly, but leveraging attention maps, [55] proposed a guided masking strategy for MAE, rather than a random policy. Periodically, all images are passed through the encoder and their latent feature representation of the [CLS] token from the last attention layer is extracted in order to compute an importance map of all patches. Using this, the masked patches are sampled (higher chance for more salient patches) and then estimated via reconstruction, while a portion of the least important patches are completely put aside (i.e. not fed to the encoder).
@@
[67] presented a self-supervised masking-based method for the multi-agent reinforcement learning (RL) setting. For a given timestamp, the observations of all agents are taken. Some of them are masked and replaced by the values of the previous timestamp, eventually being encoded in a latent space. Then, a reconstruction model is used to generate the original agents’ feature representations from the masked sequence. The authors assert that the resulting feature space is stimulated to learn more about the interaction between agents, while being more temporal aware.
@@ 3.2 Target Features
In Figure 8, we illustrate the main categories of target features employed across various MIM models. These features can be categorized in two primary groups, low-level and high-level features. The low-level category encompasses representations such as pixels [9, 2], 3D [69, 70, 71, 72], HOG [73] and spectral features [74]. Methods based on low-level features often employ high masking ratios to encourage the learning of sufficiently abstract representations that are necessary to enhance performance on downstream tasks. The methods that use high-level features as target features can be split into two groups, namely those that rely on a pre-trained teacher model [75, 76, 77, 78, 79], and those in which the teacher and student are trained simultaneously [80, 41, 23, 81, 82, 83]. In the remainder of this section, we briefly describe the MIM pipelines that primarily contribute through the types of target features they employ.
@@
A handful of studies used 3D data as target features within their proposed MIM pipeline. [69] proposed the masked image modeling pre-training for 3D meshes. Inspired by ViT, they begin by grouping together faces (each face being represented by a 10D vector) into a patch. Then, the authors adopt the transformer architecture, using the 3D location of the center of each patch to compute the positional embedding. The algorithm follows MAE: patches are randomly masked with a high masking ratio, the visible patches are passed through the encoder, then the resulting latent embeddings are concatenated with masked tokens (associated with the masked patches), and subsequently decoded. The objective is not only to reconstruct the masked patches (by predicting the coordinates of the vertices), but also the faces of the patch (the 10D representations). Aiming to learn more robust representations with 2D vision models, Mask3D [71] enhanced the 2D feature representations of the ViT backbone [22] by integrating 3D priors into the training pipeline. Mask3D utilized RGB-D data within a self-supervised framework, where both color images and corresponding dense depth maps are masked and processed through dual encoders. These encoders project the data into a higher-dimensional space, enabling a decoder to accurately reconstruct the dense depth map. This method enriches the capability of ViT to handle spatial depth alongside traditional 2D data. Focusing on point clouds, GeoMAE [72] is an adaptation of the MAE framework to point clouds. This method changed the usual reconstruction objective and replaced it with centroid prediction, normal estimation and curvature prediction. This change showed significant improvements in downstream tasks such as object detection, segmentation and multi-object tracking. In a semi-supervised setting, given a labeled (source) and an unlabeled (target) 3D point cloud dataset, the aim of [84] is to transfer the knowledge from the latter to the former by embedding information about common features in an encoder. This model is trained simultaneously on both datasets, but with different objectives. On the one hand, training on the source dataset is performed in a supervised setting on a specific task. On the other hand, points from the target dataset are randomly masked in arbitrary areas and the objective is to estimate their cardinality (number of points in the neighborhood), position and normal vectors. Therefore, the model benefits from unlabeled data by encapsulating information about the structure of objects.
@@ 3.3 Objective Function
[25] formally introduced the self-distillation masked autoencoder framework. An input image is divided into patches, some of which are randomly masked. Two encoder-decoder networks (teacher and student) are used to reconstruct the original image. A new objective is employed to minimize the distance between the predictions of the teacher and the student. While the student is trained using gradient descent, the teacher’s weights are computed as an exponential moving average of the student’s. [90] proposed a similar method based on Siamese networks. One network, which is fed with a highly masked input image, tries to predict the tokens from the other network, which is given the complete image. The work of [4] is built on the same idea, but combines the reconstruction objective with a contrastive loss. As presented in Figure 11, the authors achieve this by using two branches, one for each strategy. The first one uses an encoder and a pixel decoder, which are updated at every step. The second one employs a distinct encoder, which is updated as an exponential moving average from the other encoder. It also uses a projection layer and a feature decoder that have the same output vector space. Two views that have a different a shift are created from the input image, one being passed through the first branch (as well as being masked), while the other is fed into the second branch (unaltered). The first branch follows the original MAE framework, with the reconstruction objective. The second branch applies a contrastive loss between the projected embedded features (with the latter encoder) of the second view and the decoded embedded features (i.e. from the former encoder) of the first view. A similar hybrid approach is used by [93]. The authors apply a reconstruction loss to masked images and a contrastive loss to masked video frames, enabling their method to be effective in both image and video tasks.
@@
Figure 12: First step of AMAE, courtesy of [95] (image licensed under CC BY-SA 4.0).
Figure 13: Second step of AMAE pipeline, courtesy of [95] (image licensed under CC BY-SA 4.0). AMAE belongs to the family of reconstruction-based methods (it falls under the generic pipeline presented in Figure 2), as it employs an ${L}^{2}$ loss for masked patches reconstruction, similar to MAE.
@@ 3.4 Downstream Task
MaskGIT [97] trains a generative transformer to reconstruct randomly masked image patches. At inference time, it starts by predicting all the patches and keeps the most confident ones for the next iteration, when the rest of the patches are again masked and regenerated. The process continues for a few iterations. Overall the pipeline has actually two stages, the first one encodes the patches into visual tokens with a VQ-Encoder, and the second stage (decoder) receives masked tokens for reconstruction. Inspired by this, [7] adopted the token masking strategy in their work, but replaced the values with zero (rather than a special token) and the token embeddings are obtained with a single dense layer. Another work that studies the application of MIM in image synthesis is MDT [107]. In their work, [107] integrated latent masking in the training pipeline of a latent diffusion model.
[102] proposed to use MAE on videos. The video is split into equally-sized patches that do not overlap along any dimension (i.e. including time), as well as consisting only of two timesteps. As hypothesized and demonstrated by the authors, a high masking ratio (about 90% of the whole input video, unaware of any axis) is used due to redundant data when decoding. Positional (i.e. height and width) and temporal (i.e. time) embeddings are added to the input tokens. The architectures of both the encoder and the decoder are based on ViT. The method follows the same logic as MAE: encoding only the visible tokens, then decoding the complete set tokens. Targeting video generation, MAGVIT [101] trained a transformer-based model for conditional video generation, handling tasks like frame prediction, interpolation, and central outpainting. The model processes videos by selecting a task, and creating conditional tokens from the preprocessed video. The conditional tokens, along with masked and original tokens, form the input used to train the model, which is optimized with three objectives: refining conditional tokens, predicting masked tokens, and reconstructing original tokens. Inference is conducted autoregressively, reducing the masking ratio progressively.
@@
Several works have focused on various tasks in the medical domain. Masked Autoencoder Guided Segmentation at Pixel Resolution (MAESTER) [110] is a masked image modeling approach for the segmentation of cellular images. MAESTER incorporates the masked autoencoder in the training pipeline of a visual transformer to learn token representations relevant to the sub-cellular structure segmentation task (i.e. texture) and performs the segmentation at inference time by clustering the learned representations. [98] used masked image modeling to learn a latent space from fMRI inputs. After this stage, the authors used the learned latent representations to condition a diffusion model that is able to generate visualizations of the initial visual stimuli. [95] leveraged an MAE for detecting anomalies in chest X-rays. The first stage of their pipeline (illustrated in Figure 12) consists of initially pre-training a masked autoencoder. A classification head is then attached to the encoder (whose weights are frozen) to classify normal and abnormal samples, the latter being artificially created. During the second stage, unlabeled data is classified using the model from the previous stage, and the examples that have high confidence are kept. The last step is to employ two different copies of the pre-trained autoencoder, one for each class, and train them separately on the masked reconstruction task. At inference, multiple reconstructions are generated from both autoencoders, an anomalous prediction being detected by a large difference between the mean reconstructed images of the two modules. The steps in the second stage are presented in Figure 13. In an effort to boost the performance of computer vision models in the medical domain, [100] used visual transformers for multi-label classification of chest X-rays. The large quantity of data required for training a ViT was overcome by pre-training the model on the masked modeling task. Furthermore, the authors proved the superior performance of their method compared with CNN-based architectures. Similar to [100], [99] demonstrated the benefits of masked image modeling on 3D medical images. They adopted two strategies (original MAE and SimMIM) with various masking hyperparameters, showcasing improved results on two different segmentation tasks. [111] adopted the MAE pre-training strategy for microscopy images. Their objective was to learn a rich feature representation of the cellular morphology of the data. The extensive experiments demonstrated excellent results for predicting biological relationships. Leveraging both images and text, [112] introduced masked modeling as a pre-training strategy for medical radiography tasks by using the associated radiology reports. While the images are downsampled by half and their unmasked patches are encoded, the text reports are tokenized, randomly masked and then embedded through a look-up table. A global average pooling over the resulting visual embeddings is added to the unmasked text embeddings, and all of them are decoded to obtain the intact report. The image tokens are fed as well through another decoder in order to reconstruct the radiography at its original resolution.

### v3

@@ 2.1 Reconstruction
The first scheme that we identified revolves around the idea of masking some piece of information at any stage during the forward pass of the model, and then employing a decoder to reconstruct the missing data. We illustrate the reconstruction framework in Figure 2. Typically, the input tokens in this pipeline correspond to patches of raw pixels [9, 2]. However, there are scenarios in which these tokens reside in a latent space [32, 33]. In Figure 2, we illustrate this optional latent space projection through an image tokenizer, which can vary in complexity. For instance, [32] employed a VQ-GAN [34] tokenizer to transform images into sequences of semantic visual tokens. We emphasize that all recent advances [35, 36, 37, 38] on image tokenizers can be easily integrated into our generic formulation.
The reconstruction-based pre-training framework was concurrently introduced by [9] and [2]. Both studies employ an encoder based on the ViT architecture [22], in which a significant portion of the input tokens is masked. The encoder processes only the visible tokens. After the encoding stage, the masked tokens are replaced with a special mask token, and a decoder reconstructs their original content. In the Masked Autoencoder (MAE) proposed by [9], the decoder is lighter than the encoder, which results in an asymmetric architecture. The loss function is applied only to the output corresponding to the masked tokens.
After pre-training, the resulting encoder can be repurposed for feature extraction in downstream tasks, thanks to its strong representational capacity. Moreover, the decoder can be employed for generation, as demonstrated by [32] and [33]. The generative process involves iterative model evaluations and begins with an image in which all tokens are initially masked. During each iteration, some of the masked tokens are replaced with predictions from the model. This process continues for a fixed number of steps, with the final iteration replacing any remaining masked tokens with the model’s predictions, ultimately producing a complete image.
@@
Models: $T_{\omega}$ – the tokenizer; $E_{\theta}$ – the encoder; $D_{\phi}$ – the decoder.
@@
2:$P=T_{\omega}(P)$
3:$I_{v},I_{m}\leftarrow\mbox{mask}(P,\alpha,n)$
4:for $i\in\{1,\dots,n\}$ do
5:     if $i\in I_{v}$ then
6:         $H[i]\leftarrow E_{\theta}(p_{i})$
7:     else if $i\in I_{m}$ then
8:         $H[i]\leftarrow M$
9:$\hat{P}\leftarrow D_{\phi}(H)$
10:$\mathcal{L}\left(\phi,\varphi,\theta,M\right)\leftarrow d\left(\hat{P},P\right)$
11:$\theta\leftarrow\theta-\eta\cdot\frac{\partial\mathcal{L}}{\partial\theta}$
12:$\phi\leftarrow\phi-\eta\cdot\frac{\partial\mathcal{L}}{\partial\phi}$
@@
We formally present the reconstruction-based training strategy in Algorithm 1. The first step of the algorithm splits the input images into non-overlapping patches, resulting in a set $P$ that contains patches of the same size, namely of $h\times w$ pixels. The second step applies the tokenizer $T_{\omega}$, which can differ from one method to another. Notably, $T_{\omega}$ can be represented by the identity function when the MIM pre-training is directly applied on the raw pixels. The masking operation indicates which patches should be kept and which should be masked. The indexes of the visible and masked patches are stored in $I_{v}$ and $I_{m}$, respectively. In the next steps, the masked patches are usually dropped, and only the visible patches are processed by an encoder ($E_{\theta}$) that extracts a latent representation. Before the decoding step, the previously masked patches are replaced by a learnable representation ($M$), which resides in the latent space of the encoder. These transformations correspond to steps 3-8 in Algorithm 1.
Using the sequence formed by concatenating the masked representations with those from the encoder, the decoder ($D_{\phi}$) reconstructs the patches as depicted in step 9 of Algorithm 1. Finally, in steps 10-13 of Algorithm 1, a distance function is optimized by updating the encoder, the decoder, the projection layer, and the learnable masked representation.
@@ 2.2 Contrastive
Figure 4: A multi-level classification of Masked Image Modeling papers into various categories, based on the research directions studied in the respective papers. References are clickable links to papers. Best viewed in color.
2.3 Relation to Other SSL methods
Contrastive learning. The standard contrastive learning paradigm [42, 43, 44] is to learn representations by pulling different augmented views of the same image closer in the embedding space, while pushing views of different images apart. In Section 2.2, we show that MIM and contrastive learning are not mutually exclusive. However, if we compare the standard contrastive learning with MIM, we can identify a few different properties. First, MIM is naturally well-suited to transformer architectures, as it draws inspiration from masked language modeling, which was originally designed for such models. Second, unlike contrastive learning, MIM supports both recognition and generation tasks, as discussed in Section 2.1. In terms of representation learning, MIM encourages the model to focus on spatially meaningful features, because it reconstructs masked regions of the input. Lastly, contrastive learning relies on negative examples, which increases the batch size and often requires hard sample mining, a procedure that introduces additional computational overhead and complexity during training [45].
Non-contrastive methods. Early self-supervised learning approaches primarily focused on creating pretext tasks, e.g. solving jigsaw puzzles [13, 46], learning the arrow of time [47], or colorizing images [16]. While conceptually simple, these methods often exhibit limited generalization, because the objectives of the pretext tasks are not always aligned with those of downstream tasks. More recent studies [48, 49, 50] leverage semantic similarities between samples. The core idea is to employ a dual-network architecture, where both networks are trained to generate similar embeddings for different augmented views of the same image. Notably, this approach differs from contrastive learning, because it eliminates the need for negative examples.
Compared with MIM, the aforementioned self-supervised approaches differ in several key aspects. First, they can be less stable during training and often rely on additional techniques, such as momentum encoders, to ensure convergence. Second, unlike alternative methods, MIM inherently supports image generation. Lastly, while MIM focuses on learning spatial and contextual features through reconstruction, other self-supervised learning methods prioritize strong global representations by enforcing consistency across different views of the same image.
@@ 3.1 Masking Strategy
A number of studies used low-level information to improve the masking strategy. [68] formulated the masking strategy as as curriculum learning problem by gradually increasing the masking ratio during training. Additionally, they introduced a patch selection strategy that emphasizes masking patches with higher gradient magnitudes. The masking strategy proposed by [70] is inspired by the color noise used in image processing. Specifically, their approach introduces four filters that operate on the uniform noise typically generated for random masking. This strategy avoids the extra computational overhead often required by other informed masking methods, while still outperforming standard random masking. Rather than imposing a masking policy, several studies tried to learn the optimal masking strategy. SemMAE [58] deployed an additional stage before masking. This stage, called semantic part learning, is responsible for learning attention maps that correspond to meaningful semantic components in the image. The training of this stage is performed by embedding the class token provided by a ViT encoder into part embeddings. The resulting embeddings together with the patch embeddings provided by the same encoder are used in an attention layer. Finally, the obtained attention maps are processed by a decoder to reconstruct the original image. In the second stage, the attention maps are used for part segmentation. The masking varies from masking patches in each part to masking entire parts randomly. In a similar fashion, [57] proposed a framework in which the masking strategy is learned. A teacher network, seeing the intact image, generates a mask that is used to pre-train the student. The objective is to reconstruct the feature representation of the teacher for the masked tokens, as well as the class token [CLS] that is used in generating the mask. The teacher’s parameters are updated using an exponential moving average of the student’s weights. The experiments demonstrated the superiority of the method over other masking strategies. Instead of relying on an additional teacher network, [63] developed a masking strategy based on the direct feedback provided by the model being pre-trained. The strategy computes the reconstruction loss for each patch and then selectively masks the patches that are more challenging to reconstruct, as indicated by their higher loss values. This approach ensures that the model focuses on the most difficult aspects of the data during training. Similarly, but leveraging attention maps, [71] proposed a guided masking strategy for MAE, rather than a random policy. Periodically, all images are passed through the encoder and their latent feature representation of the [CLS] token from the last attention layer is extracted in order to compute an importance map of all patches. Using this, the masked patches are sampled (higher chance for more salient patches) and then estimated via reconstruction, while a portion of the least important patches are completely put aside (i.e. not fed to the encoder).
@@
[83] presented a self-supervised masking-based method for the multi-agent reinforcement learning (RL) setting. For a given timestamp, the observations of all agents are taken. Some of them are masked and replaced by the values of the previous timestamp, eventually being encoded in a latent space. Then, a reconstruction model is used to generate the original agents’ feature representations from the masked sequence. The authors assert that the resulting feature space is stimulated to learn more about the interaction between agents, while being more temporally aware.
@@ 3.2 Target Features
In Figure 8, we illustrate the main categories of target features employed across various MIM models. These features can be categorized in two primary groups, low-level and high-level features. The low-level category encompasses representations such as pixels [9, 2], discrete tokens [85, 86], 3D [87, 88, 89, 90], HOG [91], and spectral features [92]. Methods based on low-level features often employ high masking ratios to encourage the learning of sufficiently abstract representations that are necessary to enhance performance on downstream tasks. The methods that use high-level features as target features can be split into two groups, namely those that rely on a pre-trained teacher model [93, 94, 95, 96, 97], and those in which the teacher and student are trained simultaneously [98, 57, 23, 99, 100, 101]. In the remainder of this section, we briefly describe the MIM pipelines that primarily contribute through the types of target features they employ.
@@
A handful of studies used 3D data as target features within their proposed MIM pipeline. [87] proposed the masked image modeling pre-training for 3D meshes. Inspired by ViT, they begin by grouping faces together (each face being represented by a 10D vector) into a patch. Then, the authors adopt the transformer architecture, using the 3D location of the center of each patch to compute the positional embedding. The algorithm follows MAE: patches are randomly masked with a high masking ratio, the visible patches are passed through the encoder, then the resulting latent embeddings are concatenated with masked tokens (associated with the masked patches), and subsequently decoded. The objective is not only to reconstruct the masked patches (by predicting the coordinates of the vertices), but also the faces of the patch (the 10D representations). Aiming to learn more robust representations with 2D vision models, Mask3D [89] enhanced the 2D feature representations of the ViT backbone [22] by integrating 3D priors into the training pipeline. Mask3D utilized RGB-D data within a self-supervised framework, where both color images and corresponding dense depth maps are masked and processed through dual encoders. These encoders project the data into a higher-dimensional space, enabling a decoder to accurately reconstruct the dense depth map. This method enriches the capability of ViT to handle spatial depth alongside traditional 2D data. Focusing on point clouds, GeoMAE [90] is an adaptation of the MAE framework to point clouds. This method changed the usual reconstruction objective and replaced it with centroid prediction, normal estimation and curvature prediction. This change showed significant improvements in downstream tasks such as object detection, segmentation and multi-object tracking. In a semi-supervised setting, given a labeled (source) and an unlabeled (target) 3D point cloud dataset, the aim of [102] is to transfer the knowledge from the latter to the former by embedding information about common features in an encoder. This model is trained simultaneously on both datasets, but with different objectives. On the one hand, training on the source dataset is performed in a supervised setting on a specific task. On the other hand, points from the target dataset are randomly masked in arbitrary areas and the objective is to estimate their cardinality (number of points in the neighborhood), position and normal vectors. Therefore, the model benefits from unlabeled data by encapsulating information about the structure of objects.
@@ 3.3 Objective Function
[25] formally introduced the self-distillation masked autoencoder framework. An input image is divided into patches, some of which are randomly masked. Two encoder-decoder networks (teacher and student) are used to reconstruct the original image. A new objective is employed to minimize the distance between the predictions of the teacher and the student. While the student is trained using gradient descent, the teacher’s weights are computed as an exponential moving average of the student’s. [108] proposed a similar method based on Siamese networks. One network, which is fed with a highly masked input image, tries to predict the tokens from the other network, which is given the complete image. The work of [4] is built on the same idea, but combines the reconstruction objective with a contrastive loss. As presented in Figure 11, the authors achieve this by using two branches, one for each strategy. The first one uses an encoder and a pixel decoder, which are updated at every step. The second one employs a distinct encoder, which is updated as an exponential moving average from the other encoder. It also uses a projection layer and a feature decoder that have the same output vector space. Two views that have a different shift are created from the input image, one being passed through the first branch (as well as being masked), while the other is fed into the second branch (unaltered). The first branch follows the original MAE framework, with the reconstruction objective. The second branch applies a contrastive loss between the projected embedded features (with the latter encoder) of the second view and the decoded embedded features (i.e. from the former encoder) of the first view. A similar hybrid approach is used by [111]. The authors apply a reconstruction loss to masked images and a contrastive loss to masked video frames, enabling their method to be effective in both image and video tasks.
@@ 3.4 Downstream Task
MaskGIT [32] trains a generative transformer to reconstruct randomly masked image patches. At inference time, it starts by predicting all the patches and keeps the most confident ones for the next iteration, when the rest of the patches are again masked and regenerated. The process continues for a few iterations. Overall, the pipeline has actually two stages, the first one encodes the patches into visual tokens with a VQ-Encoder, and the second stage (decoder) receives masked tokens for reconstruction. Inspired by this, [7] adopted the token masking strategy in their work, but replaced the values with zero (rather than a special token) and the token embeddings are obtained with a single dense layer. Another work that studies the application of MIM in image synthesis is MDT [124]. In their work, [124] integrated latent masking in the training pipeline of a latent diffusion model.
[119] proposed to use MAE on videos. The video is split into equally-sized patches that do not overlap along any dimension (i.e. including time), as well as consisting only of two timesteps. As hypothesized and demonstrated by the authors, a high masking ratio (about 90% of the whole input video, unaware of any axis) is used due to redundant data when decoding. Positional (i.e. height and width) and temporal (i.e. time) embeddings are added to the input tokens. The architectures of both the encoder and the decoder are based on ViT. The method follows the same logic as MAE: encoding only the visible tokens, then decoding the complete set of tokens. Targeting video generation, MAGVIT [118] trained a transformer-based model for conditional video generation, handling tasks like frame prediction, interpolation, and central outpainting. The model processes videos by selecting a task, and creating conditional tokens from the preprocessed video. The conditional tokens, along with masked and original tokens, form the input used to train the model, which is optimized with three objectives: refining conditional tokens, predicting masked tokens, and reconstructing original tokens. Inference is conducted autoregressively, reducing the masking ratio progressively.
Figure 13: AMAE [117] belongs to the family of reconstruction-based methods (it falls under the generic pipeline presented in Figure 2), as it employs an ${L}^{2}$ loss for masked patch reconstruction, similar to MAE.
@@
Several works have focused on various tasks in the medical domain. Masked Autoencoder Guided Segmentation at Pixel Resolution (MAESTER) [127] is a masked image modeling approach for the segmentation of cellular images. MAESTER incorporates the masked autoencoder in the training pipeline of a visual transformer to learn token representations relevant to the sub-cellular structure segmentation task (i.e. texture) and performs the segmentation at inference time by clustering the learned representations. [114] used masked image modeling to learn a latent space from fMRI inputs. After this stage, the authors used the learned latent representations to condition a diffusion model that is able to generate visualizations of the initial visual stimuli. [117] leveraged an MAE for detecting anomalies in chest X-rays. The first stage of their pipeline consists of initially pre-training a masked autoencoder. A classification head is then attached to the encoder (whose weights are frozen) to classify normal and abnormal samples, the latter being artificially created. During the second stage, unlabeled data is classified using the model from the previous stage, and the examples that have high confidence are kept. The last step is to employ two different copies of the pre-trained autoencoder, one for each class, and train them separately on the masked reconstruction task. At inference, multiple reconstructions are generated from both autoencoders, an anomalous prediction being detected by a large difference between the mean reconstructed images of the two modules. The steps in the second stage are presented in Figure 13. In an effort to boost the performance of computer vision models in the medical domain, [116] used visual transformers for multi-label classification of chest X-rays. The large quantity of data required for training a ViT was overcome by pre-training the model on the masked modeling task. Furthermore, the authors proved the superior performance of their method compared with CNN-based architectures. Similar to [116], [115] demonstrated the benefits of masked image modeling on 3D medical images. They adopted two strategies (original MAE and SimMIM) with various masking hyperparameters, showcasing improved results on two different segmentation tasks. [128] adopted the MAE pre-training strategy for microscopy images. Their objective was to learn a rich feature representation of the cellular morphology of the data. The extensive experiments demonstrated excellent results for predicting biological relationships. Leveraging both images and text, [129] introduced masked modeling as a pre-training strategy for medical radiography tasks by using the associated radiology reports. While the images are downsampled by half and their unmasked patches are encoded, the text reports are tokenized, randomly masked and then embedded through a look-up table. A global average pooling over the resulting visual embeddings is added to the unmasked text embeddings, and all of them are decoded to obtain the intact report. The image tokens are fed as well through another decoder in order to reconstruct the radiography at its original resolution.